OpenStudy (anonymous):

Answer to 1.3(c-2)=2(c-6) A: 2 B:-6 C: no solution D: identity 2.0.5(m+4)=3(m-1)

OpenStudy (anonymous):

Distribute the 3 and 2. solve for Hint: there will be a solution

OpenStudy (anonymous):

solve for c*

OpenStudy (anonymous):

There are a couple of ways to do this, the easiest for this one will be to just multiply it all out on each side first. Therefore: 3(c-2)=2(c-6) is equal to 3c-6=2c-12 Get the c's on one side and solve